40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 17 03:30:13 2012, Den, Ayaka, Update, Alignment, red in arms 
    Reply  Sat Nov 17 18:02:14 2012, Den, Update, Alignment, red in arms 
       Reply  Sat Nov 17 22:50:17 2012, Koji, Update, Alignment, red in arms 
       Reply  Mon Nov 19 15:12:52 2012, Jenne, Update, Alignment, Yarm locked IR 
Message ID: 7720     Entry time: Sat Nov 17 03:30:13 2012     Reply to this: 7721
Author: Den, Ayaka 
Type: Update 
Category: Alignment 
Subject: red in arms 

We aligned accurately 00 green in yarm, changed voltage on PZT2 to see red flashing at TRY at the normalized level 0.2-0.3. The plan was to lock yarm using POY11 and green from other side, maximize red TRY by adjusting PZT2. But POY11 does not go out of the vacuum, so we adjusted TRY by flashing. 2 DOFs of PZT2 is not enough to match 4 DOFs of red beam so we adjusted both PZT2 and cavity mirrors. TRY flashing is 0.5-0.6 and green is still locking to 00 though its transmission is not maximized. We'll fix it later by adjusting input green beam.

Next we wanted to get red beam on TRX PD. Beam steering was done by BS only. We misaligned BS in pitch and excited BS angle motion by 1000 counts. We could see red beam moving on the wall of ETMX chamber. We moved it to ETMX mirror frame, estimated position of the mirror center and moved BS to this position. The beam should be approximately in the middle. For now we can not see red beam on the camera at ETMX table, more work is needed.

ELOG V3.1.3-